2.0g of a hydrated salt XnH2O was heated to a constant mass 1.09g of the anhydrous salt X was left.given that the relative formula mass of X=152 find n
1
Expert's answer
2020-04-03T11:50:50-0400

M (H2O) = 18n

m (H2O) = 2.0 - 1.09 = 0,91 g

M (X) = 152

Ratio:

1.09/152 = 0.91/18n

n = 7